# Diff-BBO: Diffusion-Based Inverse Modeling for Black-Box Optimization

### v1

@@ 3.2 Conditional Diffusion Model Training
Diffusion Models (Sohl-Dickstein et al., 2015; Song et al., 2020) are probabilistic generative models that learn distributions through an iterative denoising process. These models consist of three components: a forward diffusion process that produces a series of noisy samples by adding Gaussian noise, a reverse process to reconstruct the original data samples from the noise, and a sampling procedure to generate new data samples from the learned distribution. Let the original sample be $\mathbf{x}_{0}$ and $t$ be the diffusion step. For conditional diffusion models, a conditional variable $y$ is added to both the forward process as $q\left(\bm{x}_{t}|\bm{x}_{t-1},y\right)$ and reverse process as $p_{\theta}\left(\bm{x}_{t-1}\mid\bm{x}_{t},y\right),~{}\forall t\in[T]$.
The reverse process begins with the standard Gaussian distribution $p(\bm{x}_{T})=\mathcal{N}(\bm{0},\bm{I})$, and denoises $\bm{x}_{t}$ to recover $\bm{x}_{0}$ through the following Markov chain with reverse transitions:
During training, $\Sigma_{\theta}$ is empirically fixed, and $\mu_{\theta}$ is reparametrized by a trainable denoise function $\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t,y\right)$, which is used to estimate the noise vector $\epsilon$ that was added to input $\bm{x}_{t}$, and is trained by minimizing a reweighted version of the evidence lower bound (ELBO):
Note that the loss in Equation 2 (Ho et al., 2020) for $\bm{\epsilon}_{\theta}$ is denoising score matching for all time step $t$, which estimates the gradient of the log probability density of the noisy data (a.k.a. score function): $\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t,y\right)\approx-\sigma_{t}\nabla_{\bm{x}}\log p(\bm{x}\mid y)$. We further denote the score function as $s_{\theta}(\bm{x}_{t},y,t):=-\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t,y\right)/\sigma_{t}$.
@@
Instead of learning a fixed deterministic $\theta$ from a deterministic neural network, we are interested in learning its Bayesian posterior to further understand and improve the model’s performance as well as its reliability with uncertainty quantification. In Bayesian settings, we consider the model parameters $\theta\in\Theta$, where $\Theta$ is the parameter space, and maintain its posterior distribution $p(\theta|\mathcal{D})$, which is learned from training data $\mathcal{D}$. By choosing $\theta$ from its posterior, essentially we sample a score function $\widetilde{s}_{\theta}(\bm{x}_{t},y,t)$ from the probability distribution $p(s_{\theta}~{}|~{}\bm{x}_{t},y,t,\mathcal{D})=\mathcal{N}(s_{\theta}(\bm{x}_{t},y,t),\Sigma_{s_{\theta}}(\bm{x}_{t},y,t))$, whose expected value is $s_{\theta}(\bm{x}_{t},y,t)$, and variance is a diagonal covariance matrix $\Sigma_{s_{\theta}}(\bm{x}_{t},y,t)$.
@@
Specifically, we adopt classifier-free guidance as in (Ho and Salimans, 2022) to eliminate the requirement of training a separate classifier model. We jointly train an unconditional diffusion model $p_{\theta}(\bm{x})$ parameterized by $\epsilon_{\theta}(\mathbf{x},t,\emptyset)$ and a conditional diffusion model $p_{\theta}(\bm{x}|y)$ parameterized by $\epsilon_{\theta}(\mathbf{x},t,y)$ by minimizing the following loss function:
@@
$\mathcal{L}_{\mathrm{cdif}}=\mathbb{E}_{\bm{x}_{0},y,\bm{\epsilon},t,\bm{x}_{t},\lambda}\left[w\left(t\right)\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t,(1-\lambda)y+\lambda\emptyset\right)\right\|_{2}^{2}\right],$ (3)
@@
where $\bm{x}_{0}\sim q(\bm{x}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}),t\sim\mathcal{U}(0,T),\bm{x}_{t}\sim q(\bm{x}_{t}\mid\bm{x}_{0}),\lambda\sim\mathrm{Bernoulli}(p_{\mathrm{uncond}})$, and $p_{\mathrm{uncond}}$ is the probability of setting $y$ to the unconditional information $\emptyset$. The overall Diff-BBO framework with the conditional diffusion model included in shown in Figure 2.
@@
Figure 2: Black-box optimization framework using the conditional diffusion model as the inverse model. The overall framework includes $4$ stages. 1. Train the conditional diffusion model given the current training dataset. 2. Compute the acquisition function and select the optimal $y^{*}$ to condition on. 3. Generate samples $\{\mathbf{x_{0}}\}$ conditioned on $y^{*}$. 4. Query the oracle given generated samples $\{\mathbf{x_{0}}\}$ and update the training dataset.
@@ 4 Acquisition Function Design
In this section, we analyze the uncertainty of Diff-BBO, decomposing the uncertainty into aleatoric and epistemic uncertainty. Based on the uncertainty decomposition, we propose a novel acquisition function called Uncertainty-aware Exploration (UaE). We prove that by achieving a balance between high objective values and low epistemic uncertainty, UaE effectively provides a near-optimal solution to the online BBO problem.
@@ 4.1 Uncertainty Quantification on Conditional Diffusion Model
The optimization problem defined in Equation 1 presents a probabilistic formulation of the online BBO problem using inverse modeling. Instead of searching for a single optimal point $\bm{x}$, it aims to learn a parameterized distribution $p_{\theta}(\bm{x}\mid y,\mathcal{D})$ for a given $y$, and sampling from this predictive distribution. As such, we resort to the tools of Bayesian inference to solve this task. More specifically, given an observed value $y$ of a sample $\bm{x}$, the objective of Bayesian inference is to estimate the predictive distribution:
@@
$\widehat{{\mathbb{E}}}_{\theta}[p_{\theta}(\bm{x}\mid y)]=\frac{1}{M}\sum_{i=1}^{M}p_{\theta_{i}}(\bm{x}\mid y).$
By Equation 4, we recognize that the uncertainty arises from two sources: uncertainty in deciding parameter $\theta$ from its posterior $p(\theta|\mathcal{D})$ and uncertainty in generating sample $\bm{x}$ from a fixed diffusion model $p_{\theta}(\bm{x}\mid y)$ after $\theta$ is chosen. Here, let us first consider the problem of how to capture the uncertainty for a fixed diffusion model. In fact, the uncertainty in generating $\bm{x}$ can be explicitly traced through the denoising process. More specifically, Section 4.1 provides analytical solutions to compute the uncertainty on a single denoising process of general score-based conditional diffusional models. It provides theoretical insights of how uncertainty is being propagated through the reverse denoising process both in discrete time and continuous time, which is characterized through the lens of stochastic differential equations (SDEs) of the Ornstein–Uhlenbeck (OU) process. Detailed proofs can be found in Appendix A. {restatable}theoremuncertaintyPropagation (Uncertainty propagation) Let $t\in[T]$ be the diffusion step, $s_{\theta}(\bm{x},y,t)$ be the score function of the corresponding diffusion model $p_{\theta}(\bm{x}\mid y)$. For a single conditional diffusional model $p_{\theta}(\bm{x}\mid y)$, the uncertainty in generating a sample $\bm{x}$ can be analytically traced through the discrete-time reverse denoising process as follows:
@@
Nevertheless, performing exact Bayesian inference for uncertainty quantification when training diffusion models requires non-trivial efforts and can be computationally demanding. Hence, we introduce a practically-efficient uncertainty decomposition based on Equation 4.
@@ 4.2 Uncertainty Decomposition
To systematically analyze the effect of uncertainty in inverse modeling, we further provide a decomposition in terms of the aleatoric uncertainty and its epistemic counterpart.
The aleatoric uncertainty of the inverse problem is captured by the variance of the likelihood $p_{\theta}(\bm{x}\mid y)$, which is proportional to the variance of the measurement noise during sample generation, irreducible and task-inherent. To estimate the aleatoric uncertainty, we can Monte Carlo (MC) sample $\bm{x}$ for $N$ times from a learned likelihood function $p_{\theta}(\bm{x}\mid y)$ for fixed $y,\theta$.
In contrast, the epistemic uncertainty is captured through the variance of the posterior distribution $p(\theta\mid\mathcal{D})$, which is proportional to the variance of the score network, and is reducible with the increase of training data. Recall that $\Theta$ is the parameter space that contains all possible model parameters $\theta$, which are used to generate samples from the predictive distribution $p(\bm{x}\mid y,\mathcal{D})$. As the dataset size and quality grows, the variance of the posterior distribution shrinks, corresponding to the reduction of epistemic uncertainty in learned parameters $\theta\sim p(\theta\mid\mathcal{D})$.
To estimate the epistemic uncertainty, we use ensemble techniques. During the inference time, by initializing the trained ensemble models with different random seeds, we first sample $M$ model parameters $\{\theta_{i}\}_{i=1}^{M}$ to simulate $M$ conditional diffusion models. Then we generate $N$ samples $\{\bm{x}_{j}\}_{j=1}^{N}$ for each diffusion model with corresponding parameter $\theta_{i},~{}\forall i\in[M]$. Combining the above gives a practical way to decompose and estimate the two types of uncertainty, whish is formally described in Proposition 1.
@@
At each iteration $k\in[K]$, the overall uncertainty in inverse modeling can be decomposed into its aleatoric and epistemic components, which can be empirically measured as follows:
$\Delta_{\text{aleatoric }}(y,\mathcal{D})$ $=\mathbb{E}_{\theta_{i}\sim p(\cdot\mid\mathcal{D})}\left[\operatorname{Var}_{\bm{x}_{i,j}\sim p_{\theta_{i}}(\cdot\mid y)}\left(\|\bm{x}_{i,j}\|\right)\right],~{}~{}\forall i\in[M],j\in[N];$ (6)
$\Delta_{\text{epistemic }}(y,\mathcal{D})$ $=\operatorname{Var}_{\theta_{i}\sim p(\cdot\mid\mathcal{D})}\left(\mathbb{E}_{\bm{x}_{i,j}\sim p_{\theta_{i}}(\cdot\mid y)}\left[\|\bm{x}_{i,j}\|\right]\right),~{}~{}\forall i\in[M],j\in[N].$
4.3 Uncertainty-aware Exploration.
At each iteration $k\in[K]$ of Dif-BBO algorithm, the acquisition function proposes an optimal scalar value $y_{k}^{*}$ as follows:
$y_{k}^{*}=\operatorname{argmax}_{y}\alpha(y,\mathcal{D}),$
which is used to generate $\bm{x}$ in the design space using conditional difussion model.
Note that to design an effective acquisition function for inverse modeling, we need to achieve a balance between high objective values $y$ and low epistemic uncertainty. On the one hand, it is advantageous to focus on the regions in $\mathcal{X}$ whose corresponding $y$ is of high values. As function evaluations are expensive to perform, we prefer to generate samples $\bm{x}$ conditioned on higher $y$, and only query the oracle for such promising samples to solve the black-box optimization task. On the other hand, we employ the epistemic uncertainty to gauge the error in the trained diffusion model. Specifically, it helps reduce the approximation error between $y_{k}^{*}$ and the reconstructed function value $\max_{j\in[N]}f(\bm{x}_{j})$, where $f(\cdot)$ is the black-box oracle, and $\bm{x}_{j}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D}),\forall j\in[N]$.
We introduce the Uncertainty-aware Exploration (UaE) as our designed acquisition function:
@@
which utilizes the uncertainty estimation on conditional diffusion model as given in Proposition 1. As shown later, by balancing the exploration-exploitation trade-off, UaE provides an effective way to solve the online BBO problem.
@@ 4.4 Sub-optimality of UaE
To quantify the quality of generated samples, we theoretically analyze the sub-optimality performance gap between $y_{k}^{*}$ and reconstructed value at each iteration. In particular, Section 4.4 and Section 4.4 demonstrate that such sub-optimality gap can be effectively handled in inverse modeling, with proofs deferred to Appendix B. We first show that by using conditional diffusion model, the expected error of the sub-optimality performance gap is zero.
@@
$\Delta(p_{\theta},y_{k}^{*})=\left|y_{k}^{*}-\max_{j\in[N]}f(\bm{x}_{j})\right|,~{}~{}\mathrm{where}~{}~{}\bm{x}_{j}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D}),~{}~{}\forall j\in[N].$ (8)
Assume that there exists some $\theta^{*}\sim p(\theta|\mathcal{D})$ that produces a predictive distribution $p_{\theta^{*}}(\cdot\mid\mathcal{D})$ such that it is able to generate a sample $x^{*}$ that perfectly reconstructs $y_{k}^{*}$. It can be shown that
${\mathbb{E}}\left[{\Delta(p_{\theta},y_{k}^{*})}\right]=0,$
and the empirical estimator $\widehat{{\mathbb{E}}}\left[{\Delta(p_{\theta},y_{k}^{*})}\right]$ is unbiased.
Section 4.4 suggests that in expectation, the reconstructed function value $\max_{j\in[N]}f(\bm{x}_{j})$ is able to accurately recover the provided conditional information $y_{k}^{*}$. Hence, in order to obtain a reasonable estimator for the optimization problem, the remaining concern goes to the variance of the gap defined in Equation 8, which is further evaluated in Section 4.4.
@@
theoremyBoundVar (Sub-optimality bound) At each iteration $k\in[K]$, suppose $M$ model parameters $\{\theta_{i}\}_{i=1}^{M}$ are generated from the ensemble model for some fixed dataset $\mathcal{|D|}$. Suppose function $f$ is $L$-Lipschitz, it can be shown that the variance of the sub-optimality performance gap is equivalent to the epidemic uncertainty:
$\mathrm{Var}_{\theta_{i}\sim p(\cdot\mid\mathcal{D})}\left(\Delta(p_{\theta_{i}},y_{k}^{*})\right)\leq cL^{2}\Delta_{\mathrm{epistemic}}(y_{k}^{*},\mathcal{D}),$ (9)
where $c$ is some universal constant. Section 4.4 shows that the variance of the sub-optimality performance gap can be upper bounded by the epistemic uncertainty of diffusion model. Therefore, our proposed acquisition function achieves the balance between high objective value and low epistemic uncertainty.
Finally, we prove in Section 4.4 that by adopting UaE for inverse modeling to guide the selection of generated samples for solving BBO problems, we can obtain a near-optimal solution for the optimization problem defined in Equation 1. The proof is available in Appendix C.
@@
theoremoptiEqn Let $\mathcal{Y}$ be the constructed candidate set at each iteration $k\in[K]$ in Algorithm 1. By adopting UaE as the acquisition function to guide the sample generation process in conditional diffusion model, Diff-BBO (Algorithm 1) achieves a near-optimal solution for the online BBO problem defined in Equation 1:
$\max_{y_{k}\in\mathbb{R}}\sum_{k=1}^{K}f(\bm{x}_{k}),~{}~{}\bm{x}_{k}\sim p_{\theta}(\cdot\mid y_{k},\mathcal{D}),~{}~{}\theta\in\Theta~{}~{}\Leftrightarrow\max_{y_{k}\in\mathcal{Y}}\sum_{k=1}^{K}\alpha(y_{k},\mathcal{D}).$
As a result, equipped with the novel design of UaE, Diff-BBO is a theoretically sound approach utilizing inverse modeling to effectievely solve the online BBO problem.

### v3

@@ 4.2 Conditional Diffusion Model Training
Instead of estimating a set of deterministic parameters $\theta$ from a deterministic neural network, we are interested in learning its Bayesian posterior to further understand and improve the model’s performance as well as its reliability with uncertainty quantification. In Bayesian setting, we consider the model parameters $\theta\in\Theta$, where $\Theta$ is the parameter space, and maintain its posterior distribution $p(\theta|\mathcal{D})$, which is learned from training data $\mathcal{D}$. By choosing $\theta$ from its posterior, essentially we sample a score function $\widetilde{s}_{\theta}(\bm{x}_{t},y,t)$ from the probability distribution $p(s_{\theta}~{}|~{}\bm{x}_{t},y,t,\mathcal{D})=\mathcal{N}(s_{\theta}(\bm{x}_{t},y,t),\Sigma_{s_{\theta}}(\bm{x}_{t},y,t))$, whose expected value is $s_{\theta}(\bm{x}_{t},y,t)$, and variance is a diagonal covariance matrix $\Sigma_{s_{\theta}}(\bm{x}_{t},y,t)$.
@@
Specifically, we adopt classifier-free guidance as in (Ho and Salimans, 2022) to eliminate the requirement of training a separate classifier. We jointly train an unconditional diffusion model $p_{\theta}(\bm{x})$ parameterized by $\epsilon_{\theta}(\mathbf{x},t,\emptyset)$ and a conditional diffusion model $p_{\theta}(\bm{x}|y)$ parameterized by $\epsilon_{\theta}(\mathbf{x},t,y)$ by minimizing the following loss function:
@@
$\mathcal{L}=\mathbb{E}_{\bm{x}_{0},y,\bm{\epsilon},t,\bm{x}_{t},\lambda}\left[w\left(t\right)\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t,(1-\lambda)y+\lambda\emptyset\right)\right\|_{2}^{2}\right],$ (4)
@@
where $\bm{x}_{0}\sim q(\bm{x}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}),t\sim\mathcal{U}(0,T),\bm{x}_{t}\sim q(\bm{x}_{t}\mid\bm{x}_{0}),\lambda\sim\mathrm{Bernoulli}(p_{\mathrm{uncond}})$, and $p_{\mathrm{uncond}}$ is the probability of setting $y$ to the unconditional information $\emptyset$.
@@
Figure 1: Diffusion-BBO framework using the conditional diffusion model as the inverse surrogate model. It includes $4$ stages: 1. Train the conditional diffusion model given the training dataset. 2. Compute the acquisition function and select the optimal $y^{*}$ to condition on. 3. Generate samples $\{\mathbf{x_{0}}\}$ conditioned on $y^{*}$. 4. Query the oracle given generated samples $\{\mathbf{x_{0}}\}$ and update the training dataset.
@@ 5 Acquisition Function Design
In this section, we propose a novel acquisition function called Uncertainty-aware Exploration (UaE) for Diffusion-BBO. We first analyze the uncertainty of the conditional diffusion model from both theoretical and practical perspectives, decomposing the uncertainty into the aleatoric and epistemic components. Based on the uncertainty decomposition, we then propose the acquisition function UaE. We prove that by achieving a balance between high objective values and low epistemic uncertainty, UaE effectively provides a near-optimal solution for online BBO.
@@ 5.1 Uncertainty Quantification on Conditional Diffusion Model
The optimization problem defined in Equation 3 presents a probabilistic formulation of the online BBO problem using inverse modeling. Instead of searching for a single optimal point $\bm{x}$, it aims to learn a parameterized distribution $p_{\theta}(\bm{x}\mid y,\mathcal{D})$ for a given $y$, and sample from this predictive distribution. As such, we resort to the tools of Bayesian inference to solve this task. More specifically, given an observed value $y$ of a sample $\bm{x}$, the objective of Bayesian inference is to estimate the predictive distribution:
@@
$\widehat{{\mathbb{E}}}_{\theta}[p_{\theta}(\bm{x}\mid y)]=\frac{1}{M}\sum\nolimits_{i=1}^{M}p_{\theta_{i}}(\bm{x}\mid y).$
By Equation 5, we recognize that the uncertainty arises from two sources: uncertainty in deciding parameter $\theta$ from its posterior $p(\theta|\mathcal{D})$ and uncertainty in generating sample $\bm{x}$ from a fixed diffusion model $p_{\theta}(\bm{x}\mid y)$ after $\theta$ is chosen. Before proceeding with the uncertainty decomposition in the conditional diffusion model, it is crucial to understand how to capture the overall uncertainty when using a diffusion model to generate $\bm{x}$. Essentially, it can be explicitly traced through the denoising process. More specifically, Section 5.1 provides analytical solutions to compute the uncertainty on a single denoising process of general score-based conditional diffusional models. It offers theoretical insights of how uncertainty is being propagated through the reverse denoising process both in discrete time and continuous time, which is characterized through the lens of stochastic differential equations (SDEs) of the Ornstein–Uhlenbeck (OU) process. Detailed proofs can be found in Appendix A. {restatable}theoremuncertaintyPropagation (Uncertainty propagation) Let $t\in[T]$ be the diffusion step, $s_{\theta}(\bm{x},y,t)$ be the score function of the corresponding diffusion model $p_{\theta}(\bm{x}\mid y)$. For a single conditional diffusional model $p_{\theta}(\bm{x}\mid y)$, the uncertainty in generating a sample $\bm{x}$ can be analytically traced through the discrete-time reverse denoising process as follows:
@@
While Section 5.1 establishes the existence of closed-form solutions to quantify uncertainty based on the intrinsic properties of diffusion models, performing exact Bayesian inference when training diffusion models in practice requires non-trivial efforts and can be computationally demanding. Hence, in Section 5.2, we will introduce a practically-efficient approach to quantify and decompose the uncertainty based on Equation 5 .
@@ 5.2 Uncertainty Decomposition
To systematically analyze the effect of uncertainty in our inverse modeling approach using the conditional diffusion model, we now provide a practical method to perform uncertainty decomposition in terms of the aleatoric component and its epistemic counterpart.
The aleatoric uncertainty is captured by the variance of the likelihood $p_{\theta}(\bm{x}\mid y)$, which is proportional to the variance of the measurement noise during sample generation, irreducible and task-inherent. To estimate the aleatoric uncertainty, we can Monte Carlo (MC) sample $\bm{x}$ for $N$ times from a learned likelihood function $p_{\theta}(\bm{x}\mid y)$ for fixed $y,\theta$.
In contrast, the epistemic uncertainty is captured through the variance of the posterior distribution $p(\theta\mid\mathcal{D})$, which is proportional to the variance of the score network, and is reducible with the increase of training data. Recall that $\Theta$ is the parameter space that contains all possible model parameters $\theta$, which are used to generate samples from the predictive distribution $p(\bm{x}\mid y,\mathcal{D})$. As the dataset size and quality grows, the variance of the posterior distribution shrinks, corresponding to the reduction of epistemic uncertainty in learned parameters $\theta\sim p(\theta\mid\mathcal{D})$. To estimate the epistemic uncertainty, we use the ensemble approach. We train $M$ conditional diffusion models by initializing them with different random seeds and obtain $M$ model parameters $\{\theta_{i}\}_{i=1}^{M}$. Then we generate $N$ samples $\{\bm{x}_{j}\}_{j=1}^{N}$ for each diffusion model with corresponding parameter $\theta_{i},~{}\forall i\in[M]$. Combining the above gives a systematic way to decompose and estimate the two types of uncertainty in practice, which is formally described in Proposition 1.
@@
At each iteration $k\in[K],\forall i\in[M],j\in[N]$, the overall uncertainty in inverse modeling can be decomposed into its aleatoric and epistemic components, which can be empirically measured as follows:
$\Delta_{\text{aleatoric }}(y,\mathcal{D})$ $=\mathbb{E}_{\theta_{i}\sim p(\cdot\mid\mathcal{D})}\left[\operatorname{Var}_{\bm{x}_{i,j}\sim p_{\theta_{i}}(\cdot\mid y)}\left(\|\bm{x}_{i,j}\|\right)\right];$ (7)
$\Delta_{\text{epistemic }}(y,\mathcal{D})$ $=\operatorname{Var}_{\theta_{i}\sim p(\cdot\mid\mathcal{D})}\left(\mathbb{E}_{\bm{x}_{i,j}\sim p_{\theta_{i}}(\cdot\mid y)}\left[\|\bm{x}_{i,j}\|\right]\right).$
5.3 Uncertainty-aware Exploration
At each iteration $k\in[K]$ of Diffusion-BBO algorithm, the acquisition function $\alpha(y_{k},\mathcal{D})$ proposes an optimal scalar value $y_{k}^{*}$ as follows:
$y_{k}^{*}=\operatorname{argmax}_{y_{k}}\alpha(y_{k},\mathcal{D}),$
which is used to generate $\bm{x}$ in the design space using conditional diffussion model.
Intuitively, designing an effective acquisition function for inverse modeling requires a balance between high objective values $y_{k}$ and low epistemic uncertainty of $y_{k}$. On the one hand, it is advantageous to generate samples $\bm{x}$ conditioned on higher $y_{k}$ and query the oracle with these promising samples for online BBO. On the other hand, the epistemic uncertainty of $y_{k}$ can be used to gauge the approximation error between $y_{k}^{*}$ and the reconstructed function value $\max_{j\in[N]}f(\bm{x}_{j})$, where $f(\cdot)$ is the black-box oracle, and $\bm{x}_{j}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D}),\forall j\in[N]$.
Therefore, we propose the acquisition function named Uncertainty-aware Exploration (UaE):
@@ 5.3 Uncertainty-aware Exploration
which utilizes the uncertainty estimation on conditional diffusion model as given in Proposition 1. It effectively penalizes the candidates for which the model is less certain. As shown later, by balancing the exploration-exploitation trade-off, UaE provides an effective way to solve the online BBO problem.
@@ 5.4 Sub-optimality of UaE
To evaluate the quality of generated samples, we theoretically analyze the sub-optimality performance gap between $y_{k}^{*}$ and reconstructed value at each iteration. In particular, Section 5.4 and Section 5.4 demonstrate that such sub-optimality gap can be effectively handled in inverse modeling, contributing to the optimality of our proposed algorithm as established in Section 5.4. We emphasize that the purpose of Section 5.4 and Section 5.4 is to support Section 5.4 by validating the design of UaE. These results are designed to hold universally with probability one, independent of hyperparameters such as batch size. As a result, the bounds presented are intentionally conservative and are not tightened to rely on high-probability guarantees. Detailed proofs are deferred to Appendix B.
We first show that by using conditional diffusion model, the expected error of the sub-optimality performance gap can be effectively bounded under mild assumptions.
@@
$\Delta(p_{\theta},y_{k}^{*})=\left|y_{k}^{*}-\max_{j\in[N]}f(\bm{x}_{j})\right|,~{}~{}\mathrm{where}~{}~{}\bm{x}_{j}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D}).$ (9)
Assume that there exists some $\theta^{*}\sim p(\theta|\mathcal{D})$ that produces a probability distribution $p_{\theta^{*}}(\cdot\mid\mathcal{D})$ such that it is able to generate a sample $\bm{x}^{*}$ that perfectly reconstructs $y_{k}^{*}$. Suppose function $f$ is $L$-Lipschitz and each sample is $\sigma$-subGaussian, it can be shown that
${\mathbb{E}}\left[{\Delta(p_{\theta},y_{k}^{*})}\right]\leq c_{1}L\sqrt{d}\sigma,$
where $d$ is the dimensionality of the design space, $c_{1}$ is some universal constant.
Section 5.4 suggests that in expectation, the reconstructed function value closely approximates the provided conditional information $y_{k}^{*}$, implying Diffusion-BBO is effective in searching for promising samples in the design space by utilizing the information from the objective space. Hence, to achieve a robust estimator for the online BBO problem, the primary concern shifts to controlling the variance of the sub-optimality gap defined in Equation 9, which is further assessed and evaluated in Section 5.4.
@@
theoremyBoundVar (Sub-optimality bound) At each iteration $k\in[K]$, suppose $M$ model parameters $\{\theta_{i}\}_{i=1}^{M}$ are generated from the ensemble model for some fixed dataset $\mathcal{D}$. Suppose function $f$ is $L$-Lipschitz, it can be shown that the variance of the sub-optimality performance gap of each model is bounded by the epidemic uncertainty:
$\mathrm{Var}\left(\Delta(p_{\theta_{i}},y_{k}^{*})\right)\leq c_{2}L^{2}d\sigma^{2}+c_{2}L^{2}\Delta_{\mathrm{epistemic}}(y_{k}^{*},\mathcal{D}),$ (10)
where $c_{2}$ is some universal positive constant. Section 5.4 shows that the variance of the sub-optimality performance gap is upper bounded by the epistemic uncertainty of diffusion model with some global constants. It implies that decreasing the epistemic uncertainty reduces the variance of the performance gap, leading to more reliable results. Therefore, it is crucial to achieve an effective balance between maximizing the objective value and minimizing the corresponding epistemic uncertainty when designing the acquisition function. By doing so, UaE not only explores the objective space with high-value solutions, but also ensures stability and consistency of the optimization process.
Finally, we prove in Section 5.4 that by adopting UaE for inverse modeling to guide the selection of generated samples for solving BBO problems, we can obtain a near-optimal solution for the online optimization problem defined in Equation 3. The proof is available in Appendix C.
@@
theoremoptiEqn Let $\mathcal{Y}$ be the constructed candidate set at each iteration $k\in[K]$ in Algorithm 1. By adopting UaE as the acquisition function to guide the sample generation process in conditional diffusion model, Diffusion-BBO (Algorithm 1) achieves a near-optimal solution for the online BBO problem defined in Equation 3:
$\qquad\max_{y_{k}\in\mathbb{R}}\sum\nolimits_{k=1}^{K}f(\bm{x}_{k}),~{}~{}\bm{x}_{k}\sim p_{\theta}(\cdot\mid y_{k},\mathcal{D}),~{}~{}\theta\in\Theta~{}~{}\Rightarrow\max_{y_{k}\in\mathcal{Y}}\sum\nolimits_{k=1}^{K}\alpha(y_{k},\mathcal{D}).$
Hence, equipped with the novel design of UaE, Diffusion-BBO is a theoretically sound approach utilizing inverse modeling to effectively solve the online BBO problem.